Can I please receive assistance with the sections highlighted in Yellow. Please provide an explanation as to how to calculate these figures, I am not familiar with how to calculate the Break-even, and I want to learn. Thank you!

.Year 0 Year 1 Year 2 Year 3 Totals $0 Net Economic Benefit Discount Rate (5%) PV of Benefits NPV of Benefits $40,000 0.9524 $
Answer #1

Yearly NPV cash flow = PV of benefits - PV of Costs

Overall = Cumulative NPV

Hence, Yearly NPV Cash flows are as follows

Year 0 =-30,000

Year 1 = 38096-14286 =$23810

Year 2 = 40815 - 13605 =$27210

Year 3 = 38871 - 12957 =$25,914

Overall is as follows

Year0 =-30,000

Year 1 = -30000+23810

=-$6190

Year 2 = -6190+27210

=$21020

Year 3 = 21020+25914 =$46934

Break even occurs when overall NPV cash flow reaches 0

I.e. between years 1 and 2